OpenStudy (anonymous):

According to the 2000 Census the population of Canton was 7,720. Holly Spring's population was 3,200. Canton's population increases at a rate of 80 people a year and Holly Spring's population increases by 120 people a year. Which equation may be used to determine the number of years before the populations are equal?

OpenStudy (anonymous):

7720 + 80x = 3200 + 120 x Let x= # of years. Are you able to do the rest by just solving for x?